A hiker starts at his camp and moves the following distances while exploring his surroundings: 255 m east, 125 m at an angle 30.0° north of east, and 178 m at an angle 12.0° south of east. Find his resultant displacement from camp.


1
Expert's answer
2021-10-27T08:49:15-0400

Find the eastward and southward components of displacement:

"A_N=0,\\
A_E=255.\\
B_N=125\sin30°,\\
B_E=125\cos30°,\\
C_N=-178\sin12°,\\
C_E=178\cos12°.\\\space\\
R_N=A_N+B_N+C_N=25.5,\\
R_E=A_E+B_E+C_E=537,\\\space\\
R=\sqrt{R_N^2+R_E^2}=538\text{ m},\\\space\\
\theta=\arctan\frac{R_N}{R_E}=2.7°."